OpenStudy (anonymous):

Suppose you select 2 letters at random from the word compute without replacement. Find each probability. P(1 vowel and 1 consonant)

OpenStudy (anonymous):

Note that there is 3 vowels and 4 consonants. This is a conditional probability, i.e. the probability of picking a consonant is dependent on that of picking a vowel. So we evaluate as:\[\bf P(1 \ vowel \ and \ 1 \ consonant)=P(1 \ vowel)*P(1 \ consonant)\]\[\bf = P(1 \ vowel) * P(1 \ consonant | 1 \ vowel)=\frac{ 3 }{ 7 }*\frac{ 4 }{ 6 }\]And you can finish evaluating that. Also notice that it doesn't matter whether you pick the vowel or the consonant first. The final probability of both events occuring at the end comes out the same. @baseballer20142

OpenStudy (anonymous):

@baseballer20142 Do you understand?

OpenStudy (anonymous):

yep thanks